# Efficient Diffusion Models: A Survey

### v1

@@ 4.1 Hardware-Software Co-Design
The co-design of hardware and software is pivotal for achieving efficient deployment of diffusion models in real-time and resource-constrained environments. Following algorithm-level optimizations, system-level techniques focus on integrating hardware-specific features, distributed computation, and caching mechanisms. These strategies aim to address the computational complexity and memory demands of large-scale diffusion models, enabling more practical applications across various platforms like GPUs, FPGAs, and mobile devices. One significant contribution is the work by Chen et al. (2023c), which explores GPU-aware optimizations for accelerating diffusion models directly on mobile devices. Implementing specialized kernels and optimized softmax operations reduces inference latency, achieving near real-time performance on mobile GPUs. In a related effort,  Yang et al. (2023a) propose SDA, a low-bit stable diffusion accelerator designed specifically for edge FPGAs. Utilizing quantization-aware training and a hybrid systolic array architecture as illustrated in Figure 29, SDA effectively balances computational efficiency with flexibility, handling both convolutional and attention operations efficiently. Through a two-level pipelining structure, the nonlinear operators are efficiently integrated with the hybridSA, enabling coordinated operation that enhances processing speed while reducing resource usage. Finally, SDA achieves a speedup of 97.3x when compared to ARM Cortex-A53 CPU.
Furthermore, Choi et al. (2024a) introduces a stable diffusion processor optimized for mobile platforms through patch similarity-based sparsity, mixed-precision strategies and and a Dual-mode Bit-Slice Core (DBSC) architecture that supports mixed-precision computation, which particularly targeting resource-constrained devices such as mobile platforms. Together, these optimizations significantly improve throughput and energy efficiency, making Stable Diffusion more viable for energy-sensitive applications.
@@ 4.2 Parallel Computing
Parallel computing plays a critical role in the efficient execution of diffusion models, especially given the computation-intensive nature of these algorithms in Figure 30. Recent advances in parallel computing strategies have enabled significant improvements in inference speed and scalability, often without compromising the quality of the generated output (Li et al., 2024a; Wang et al., 2024a; Li et al., 2024b; Tian et al., 2024). This section highlights several notable contributions that tackle the challenge of parallelizing diffusion models across multiple GPUs and other distributed architectures. Li et al. (2024a) introduced DistriFusion, a framework designed for distributed parallel inference tailored to high-resolution diffusion models such as SDXL. Their approach involves partitioning the model inputs into distinct patches, which are then processed independently across multiple GPUs. This method leverages the available hardware resources more effectively, achieving a 6.1x speedup on 8xA100 GPUs compared to single-card operation, all while maintaining output quality.
To address potential issues arising from the loss of inter-patch interaction, which could compromise global consistency, DistriFusion employs dynamic synchronization of activation displacements, striking a balance between preserving coherence and minimizing communication overhead.
@@
Building on the insights gained from DistriFusion, Wang et al. (2024a) further refined the distributed inference paradigm with PipeFusion. This system not only splits images into patches but also distributes the network layers across different devices, thereby reducing the associated communication costs and enabling the use of PCIe-linked GPUs instead of NVLink-connected ones. PipeFusion integrates sequence parallelism, tensor parallelism, displaced patch parallelism, and displaced patch pipeline parallelism, optimizing workflow for a wider range of hardware configurations.
@@
For applications involving add-on modules such as ControlNet and LoRA, Li et al. (2024b) developed SwiftDiffusion, as illustrated in Figure 31. This framework optimizes the serving workflow of these modules, allowing them to run in parallel on multiple GPUs. As a result, SwiftDiffusion delivers a 5x reduction in inference latency and a 2x improvement in throughput, ensuring that enhanced speed does not come at the expense of output quality.
Lastly, Tian et al. (2024) focused on the training phase with DiffusionPipe, demonstrating that pipeline parallelism can produce a 1.41x training speedup, while data parallelism contributes an additional 1.28x acceleration. Although the optimization methods for DiffusionPipe were not detailed in the notes, the combination of these parallelization strategies offers a promising direction to improve the efficiency of both the training and inference pipelines for diffusion models.
@@ 4.3 Caching Technique
Agarwal et al. (2024) proposed NIRVANA, a novel system designed to enhance the efficiency of text-to-image generation using diffusion models. Specifically, the key innovation lies in its approximate caching technique, which reduces computational costs and latency by reusing intermediate noise states from previous image generation processes. Instead of starting from scratch with every new text prompt, NIRVANA retrieves and reconditions these cached states, allowing it to skip several initial denoising steps.
Additionally, the system uses a custom cache management policy called Least Computationally Beneficial and Frequently Used (LCBFU), which optimizes the storage and reuse of cached states to maximize computational efficiency. This makes NIRVANA particularly suited for large-scale, production-level deployments of text-to-image diffusion models.
From another perspective, Ma et al. (2024b) introduces an innovative approach called DeepCache, designed to accelerate the image generation process by leveraging the temporal redundancy in the denoising steps of diffusion models, without the need for additional model training, as illustrated in Figure 33. The key insight is the observation that high-level features, such as the main structure and shape of an image, exhibit minimal changes between adjacent denoising steps. These features can be cached and reused in subsequent steps, thereby avoiding redundant computations. This method takes advantage of the U-Net architecture by combining these cached high-level features with low-level features, updating only the low-level features to reduce computational load, leading to a significant acceleration in the overall process.  Wimbauer et al. (2024) proposed Block Caching, a technique that identifies and caches redundant computations within the model’s layers during the denoising process. By reusing these cached outputs in subsequent timesteps, the method significantly speeds up inference while maintaining image quality. To optimize this caching process, they introduce an Automatic Cache Scheduling mechanism, which dynamically determines when and where to cache based on the relative changes in layer outputs over time. Additionally, the paper addresses potential misalignment issues from aggressive caching by implementing a Scale-Shift Adjustment mechanism, which fine-tunes cached outputs to align with the model’s expectations, thereby preventing visual artifacts.
@@
Recently, the application of diffusion with transformer models has yielded considerable success. Ma et al. (2024a) is concerned with the introduction of a layer caching mechanism, designated Learning-to-Cache (L2C), to accelerate diffusion transformer models. L2C exploits the redundancy between layers within the transformer architecture, dynamically caching computations from certain layers to reduce redundant calculations and lower inference costs. The implementation entails transforming the layer selection problem into a differentiable optimization problem, using interpolation to determine whether to perform a full computation or utilize cached results at different timesteps during inference. In contrast to the emphasis on layer caching,  Selvaraju et al. (2024) proposed Fast-Forward Caching (FORA), a technique designed to accelerate Diffusion Transformers (DiT) by reducing redundant computations during the inference phase. The key insight behind FORA is the observation that the outputs from the self-attention and MLP layers in a Transformer exhibit high similarity across consecutive time steps in the diffusion process. To leverage this, FORA implements a static caching mechanism where these layer outputs are cached at regular intervals, which are determined by $N$, and reused for a set number of subsequent steps, thereby avoiding recomputing similar outputs.
@@ 5 Frameworks
Diffusion model frameworks can generally be categorized based on their support for tasks such as training, fine-tuning, and inference. Specifically, frameworks that support training and/or fine-tuning are designed to provide scalable, efficient, and flexible infrastructures that enhance computational efficiency, reduce memory usage, and ensure the reliability of the training and fine-tuning processes. On the other hand, frameworks focused on inference aim to optimize throughput and reduce latency, addressing the practical needs of real-world applications. These frameworks offer various deployment options to handle diverse diffusion model tasks. Table 2 provides a summary of existing diffusion model frameworks along with their key features.
Table 2: Comparison of diffusion model frameworks.
Framework Training Fine-Tuning Inference Key Features
Diffusers Multi-model, versatile, open-source.
DALL-E × × Closed-source, API, high-quality.
OneDiff × × Optimized loading, CLI support.
LiteGen Efficient training, multi-GPU.
InvokeAI × Stable Diffusion, CLI, API.
ComfyUI-Docker × Modular, multiple models.
Grate × Image grids, model merging.
Versatile Diffusion Multi-modal, modular design.
UniDiffuser Unified, multi-modal diffusion.
Diffusers is an open-source framework by Hugging Face that provides a versatile toolset for working with diffusion models. It supports multi-model usage for text-to-image and image editing tasks. By leveraging a variety of pre-trained models and pipelines, it enables quick experimentation and fine-tuning for both research and production environments. The framework is designed to be extensible, supporting efficient model training, fine-tuning, and inference.
DALL-E is a closed-source text-to-image model developed by OpenAI. The framework is accessible via the OpenAI API, enabling users to generate high-quality images based on textual descriptions. Although it does not support direct training or fine-tuning, it provides inference capabilities that deliver state-of-the-art image synthesis. Its commercial use is constrained by the terms of the OpenAI API.
OneDiff focuses on optimizing model loading and inference processes, particularly in environments with limited computational resources. It is designed to offer fast execution and support for command-line interface (CLI) operations. OneDiff provides a way to quickly perform inference using various diffusion models, making it suitable for both developers and researchers who need a streamlined workflow.
LiteGen is a lightweight and efficient training framework tailored for diffusion models. It offers multi-GPU support and optimizes various tasks, enabling users to fine-tune pre-trained models with low computational overhead. LiteGen is especially useful for multi-task scenarios, where training speed and resource efficiency are critical.
InvokeAI is designed to facilitate both fine-tuning and inference processes. It includes support for multiple diffusion models and provides both a command-line interface (CLI) and an API for ease of use. InvokeAI is aimed at users who need a flexible and accessible toolkit for experimenting with image generation.
ComfyUI-Docker is a modular framework that supports multiple diffusion models like Stable Diffusion and ControlNet. While it offers a graphical user interface through Docker, it also allows for command-line operations. The modular design enables users to integrate various models and pipelines, providing a versatile platform for different image generation tasks.
Grate is a diffusion model toolkit specialized in generating image grids using multiple Stable Diffusion models. It supports model merging and fine-tuning to achieve varied artistic effects. Geared towards both artists and researchers, Grate offers flexibility in combining models to explore diverse image generation possibilities.
Versatile Diffusion is a framework that emphasizes multi-modal tasks and modular design. It supports training, fine-tuning, and inference, allowing users to build customized workflows tailored to specific domains. Its extensibility makes it suitable for a wide range of applications, from text-to-image generation to image-to-image translation.
UniDiffuser provides a unified framework for multi-modal diffusion, enabling text-to-image, image-to-text, and other tasks in a seamless manner. It supports training, fine-tuning, and inference, with a focus on bridging the gap between different types of data modalities. UniDiffuser is ideal for researchers who need a flexible model to handle complex diffusion tasks across various input forms.

### v2

@@ 4.1 Hardware-Software Co-Design
The co-design of hardware and software is pivotal for achieving efficient deployment of diffusion models in real-time and resource-constrained environments. Following algorithm-level optimizations, system-level techniques focus on integrating hardware-specific features, distributed computation, and caching mechanisms. These strategies aim to address the computational complexity and memory demands of large-scale diffusion models, enabling more practical applications across various platforms like GPUs, FPGAs, and mobile devices. One significant contribution is the work by Chen et al. (2023d), which explores GPU-aware optimizations for accelerating diffusion models directly on mobile devices. Implementing specialized kernels and optimized softmax operations reduces inference latency, achieving near real-time performance on mobile GPUs.
In a related effort,  Yang et al. (2023a) propose SDA, a low-bit stable diffusion accelerator designed specifically for edge FPGAs. Utilizing quantization-aware training and a hybrid systolic array architecture as illustrated in Figure 29, SDA effectively balances computational efficiency with flexibility, handling both convolutional and attention operations efficiently. Through a two-level pipelining structure, the nonlinear operators are efficiently integrated with the hybridSA, enabling coordinated operation that enhances processing speed while reducing resource usage. Finally, SDA achieves a speedup of 97.3x when compared to ARM Cortex-A53 CPU. Furthermore, Choi et al. (2024a) introduces a stable diffusion processor optimized for mobile platforms through patch similarity-based sparsity, mixed-precision strategies and and a Dual-mode Bit-Slice Core (DBSC) architecture that supports mixed-precision computation, which particularly targeting resource-constrained devices such as mobile platforms. Together, these optimizations significantly improve throughput and energy efficiency, making Stable Diffusion more viable for energy-sensitive applications.
For GPU-accelerated mobile applications requiring real-time interactivity,  Chen et al. (2023d) provides kernel-level optimizations and mobile-specific operator tuning with strongest latency. When targeting embedded edge devices with strict power budgets and static workloads,  Yang et al. (2023a) achieves superior performance via hardware-algorithm synergy. Meanwhile,  Choi et al. (2024a) firstly approach dynamic energy constraints in always-on scenarios through runtime sparsity adaptation.
@@ 4.2 Parallel Computing
Parallel computing Figure 30 plays a critical role in the efficient execution of diffusion models, especially given the computation-intensive nature of these algorithms. Recent advances in parallel computing strategies have enabled significant improvements in inference speed and scalability, often without compromising the quality of the generated output (Li et al., 2024a; Wang et al., 2024a; Li et al., 2024b; Tian et al., 2024). This section highlights several notable contributions that tackle the challenge of parallelizing diffusion models across multiple GPUs and other distributed architectures.
Li et al. (2024a) introduced DistriFusion, a framework designed for distributed parallel inference tailored to high-resolution diffusion models such as SDXL. Their approach involves partitioning the model inputs into distinct patches, which are then processed independently across multiple GPUs. This method leverages the available hardware resources more effectively, achieving a 6.1x speedup on 8xA100 GPUs compared to single-card operation, all while maintaining output quality. To address potential issues arising from the loss of inter-patch interaction, which could compromise global consistency, DistriFusion employs dynamic synchronization of activation displacements, striking a balance between preserving coherence and minimizing communication overhead.
@@
Building on the insights gained from DistriFusion, Wang et al. (2024a) further refined the distributed inference paradigm with PipeFusion. This system not only splits images into patches but also distributes the network layers across different devices, thereby reducing the associated communication costs and enabling the use of PCIe-linked GPUs instead of NVLink-connected ones. PipeFusion integrates sequence parallelism, tensor parallelism, displaced patch parallelism, and displaced patch pipeline parallelism, optimizing workflow for a wider range of hardware configurations. For applications involving add-on modules such as ControlNet and LoRA, Li et al. (2024b) developed SwiftDiffusion, as illustrated in Figure 31. This framework optimizes the serving workflow of these modules, allowing them to run in parallel on multiple GPUs. As a result, SwiftDiffusion delivers a 5x reduction in inference latency and a 2x improvement in throughput, ensuring that enhanced speed does not come at the expense of output quality. Lastly, Tian et al. (2024) focused on the training phase with DiffusionPipe, demonstrating that pipeline parallelism can produce a 1.41x training speedup, while data parallelism contributes an additional 1.28x acceleration. Although the optimization methods for DiffusionPipe were not detailed in the notes, the combination of these parallelization strategies offers a promising direction to improve the efficiency of both the training and inference pipelines for diffusion models. These methodologies encompass a hierarchical optimization, progressing from coarse-grained spatial partitioning to fine-grained module-specific adaptations, thereby empowering practitioners to strategically align parallelism configurations with underlying hardware constraints.
@@ 4.3 Caching Technique
Agarwal et al. (2024) proposed NIRVANA, a novel system designed to enhance the efficiency of text-to-image generation using diffusion models. Specifically, the key innovation lies in its approximate caching technique, which reduces computational costs and latency by reusing intermediate noise states from previous image generation processes. Instead of starting from scratch with every new text prompt, NIRVANA retrieves and reconditions these cached states, allowing it to skip several initial denoising steps. Additionally, the system uses a custom cache management policy called Least Computationally Beneficial and Frequently Used (LCBFU), which optimizes the storage and reuse of cached states to maximize computational efficiency. This makes NIRVANA particularly suited for large-scale, production-level deployments of text-to-image diffusion models. From another perspective, Ma et al. (2024b) introduces an innovative approach called DeepCache, designed to accelerate the image generation process by leveraging the temporal redundancy in the denoising steps of diffusion models, without the need for additional model training, as illustrated in Figure 33. The key insight is the observation that high-level features, such as the main structure and shape of an image, exhibit minimal changes between adjacent denoising steps. These features can be cached and reused in subsequent steps, thereby avoiding redundant computations. This method takes advantage of the U-Net architecture by combining these cached high-level features with low-level features, updating only the low-level features to reduce computational load, leading to a significant acceleration in the overall process.  Wimbauer et al. (2024) proposed Block Caching, a technique that identifies and caches redundant computations within the model’s layers during the denoising process. By reusing these cached outputs in subsequent timesteps, the method significantly speeds up inference while maintaining image quality. To optimize this caching process, they introduce an Automatic Cache Scheduling mechanism, which dynamically determines when and where to cache based on the relative changes in layer outputs over time. Additionally, the paper addresses potential misalignment issues from aggressive caching by implementing a Scale-Shift Adjustment mechanism, which fine-tunes cached outputs to align with the model’s expectations, thereby preventing visual artifacts.
@@
Recently, the application of diffusion with transformer models has yielded considerable success.  Ma et al. (2024a) is concerned with the introduction of a layer caching mechanissm, designated Learning-to-Cache (L2C), to accelerate diffusion transformer models. L2C exploits the redundancy between layers within the transformer architecture, dynamically caching computations from certain layers to reduce redundant calculations and lower inference costs. The implementation entails transforming the layer selection problem into a differentiable optimization problem, using interpolation to determine whether to perform a full computation or utilize cached results at different timesteps during inference. In contrast to the emphasis on layer caching,  Selvaraju et al. (2024) proposed Fast-Forward Caching (FORA), a technique designed to accelerate Diffusion Transformers (DiT) by reducing redundant computations during the inference phase. The key insight behind FORA is the observation that the outputs from the self-attention and MLP layers in a Transformer exhibit high similarity across consecutive time steps in the diffusion process. To leverage this, FORA implements a static caching mechanism where these layer outputs are cached at regular intervals, which are determined by $N$, and reused for a set number of subsequent steps, thereby avoiding recomputing similar outputs.
Shen et al. introduces MD-DiT, a unified framework for efficient diffusion transformers by integrating block skipping and caching strategies, enabling dynamic depth adjustment across timesteps without additional training. It optimizes computation by caching incremental changes from previous timesteps and selectively skipping non-critical blocks.  Chen et al. proposes $\Delta$-DiT, a groundbreaking framework for optimizing diffusion transformers through two synergistic innovations: step-sensitive block caching and intelligently allocated computational resources. By exploiting the temporal consistency inherent in diffusion trajectories, the system strategically reuses intermediate residual features across sequential steps. Notably, it harmonizes both computation-skipping and feature-caching mechanisms within a cohesive optimization architecture.
Generally, these methods demonstrate an trajectory from coarse temporal reuse to fine-grained layer adaptation, guiding practitioners to employ noise-state caching for prompt-variant scenarios, feature-level caching for structural consistency, and hybrid static-dynamic strategies for transformer-based models.
@@ 5 Frameworks
Table 2: Comparison of Optimization Support in Selected Diffusion Model Frameworks and Techniques. In this table, ‘Training’ indicates whether the framework accelerates the training process of diffusion models, while ‘Inference’ indicates whether it accelerates the generation process.
Framework Training Inference Key Features
[…]
Frameworks in the efficient diffusion landscape refer to high-performance tools and libraries designed to optimize training and/or inference. DM frameworks can be in general grouped based on whether they support the tasks of training and inference. Specifically, frameworks that support training aim to provide scalable, efficient, and flexible infrastructure that improves computation efficiency, reduces memory footprint, optimizes communication efficiency, and ensures reliability of the training process. Frameworks that support inference focus on optimizing inference throughput and reducing memory footprint and latency.
A subset of frameworks, including Flash Attention (Dao et al., 2022), xFormers (Lefaudeux et al., 2022), DeepSpeed (Rasley et al., 2020), and OneFlow (Yuan et al., 2021), distinguishes itself by offering comprehensive support for both training and inference phases of diffusion models. These tools, rooted in broader deep learning optimization efforts, cater to the full lifecycle of model development. Flash Attention accelerates both training and inference by optimizing attention computation, reducing memory usage and latency through techniques like tiling and recomputation, making it particularly effective for Diffusion Transformers (DiT). xFormers provides memory-efficient Transformer optimizations with a modular design, enabling flexible and efficient computation across a range of resource-intensive tasks. DeepSpeed, originally designed for large-scale model training, extends its distributed capabilities to inference, enabling scalable deployment across multiple GPUs. Similarly, OneFlow leverages its compiler-driven architecture to streamline both training and inference workflows, appealing to researchers and practitioners seeking end-to-end optimization. Together, these frameworks provide robust foundations for advancing diffusion model research and deployment, balancing flexibility with high performance.
In contrast, a growing array of tools—Stable-Fast (chengzeyi, 2024), Onediff (Contributors, 2022), DeepCache (Ma et al., 2024b), TGATE (Zhang et al., 2024b; Liu et al., ), and xDiT (Fang et al., 2024a; b; Fang & Zhao, 2024; Sun et al., 2024)—focuses exclusively on accelerating the inference stage of diffusion models, addressing the demand for rapid and resource-efficient generation. Stable-Fast optimizes the Hugging Face Diffusers ecosystem, prioritizing low-latency inference for real-time applications. Onediff builds on this trend by integrating cutting-edge techniques like caching and quantization, tailoring its acceleration to diffusion-specific workloads. DeepCache and TGATE introduce innovative caching strategies, exploiting temporal redundancies to reduce computational overhead in inference, particularly for U-Net-based models. Meanwhile, xDiT targets the emerging Diffusion Transformer (DiT) architecture, employing parallelization to enhance inference scalability. These inference-centric tools reflect a shift toward specialized optimizations, catering to the practical needs of deployment in constrained environments or novel model paradigms.
6 Future Work
Despite the significant progress made in efficient diffusion models, several promising research directions remain open. In particular, we identify the following key areas for future work:
Hybridizing Diffusion and Autoregressive Models. One promissing future direction is to explore hybridizing autoregressive and diffusion models to combine the strengths of both paradigms. This integration enables the use of key-value (KV) caching, a technique from autoregressive transformers, to accelerate diffusion and support streamable generation. A representative work is Block Diffusion(Arriola et al., 2025), which segments generation into blocks and applies autoregressive-style caching across denoising steps. This approach opens up promising opportunities to unify architectural benefits, allowing diffusion models to inherit the efficiency and online capabilities of autoregressive methods.
Without Classifier-Free Guidance (CFG). Although classifier-free guidance has been widely adopted to enhance generation quality, it introduces substantial computational overhead. Tang et al. (Tang et al., 2025) propose Model-guidance (MG), a novel training objective that eliminates the need for CFG by directly incorporating the posterior probability of conditions rather than solely modeling data distribution. This approach not only doubles inference speed by avoiding the second network forward pass required by CFG but also significantly accelerates model training with 6.5× faster convergence and approximately 60% performance improvement. Notably, when compared to concurrent methods, MG achieves state-of-the-art performance on ImageNet 256 with an FID of 1.34 while requiring only about 12% of the computational resources of comparable approaches. The effectiveness of MG suggests a promising direction for future research to explore alternative guidance mechanisms or fundamentally rethink the training paradigm to better balance computational efficiency and generation quality.
Efficient Attention Mechanisms for Video Diffusion. Bidirectional attention has become a dominant component in diffusion architectures, but it introduces substantial computational overhead that scales quadratically with sequence length. This is particularly problematic for video diffusion models where the sequence length grows linearly with the number of frames. Recent works have begun to explore efficient attention mechanisms specifically designed for diffusion transformers (DiTs). Xia et al. (Xia et al., 2025) propose AdaSpa, which leverages the hierarchical sparsity inherent in DiTs through a blockified pattern approach and adaptive search methods. Their work demonstrates that sparse characteristics of DiTs exhibit hierarchical structures between different modalities and remain invariant across denoising steps, enabling significant computational savings. Similarly, Ding et al. (Ding et al., 2025) identify tile-style repetitive patterns in 3D attention maps for video data, introducing sparse 3D attention with linear complexity relative to frame count. Their approach combines efficient attention with consistency distillation techniques to enable up to 7.8× faster generation for high-resolution videos. These developments suggest promising directions for mitigating the computational burden of attention in diffusion models without sacrificing generation quality.